OpenStudy (anonymous):

Find the exact values of the six trigonometric functions of the angle theta. Sin(theta)= 3/5, pi/2 <= theta <= pi

OpenStudy (anonymous):

\[\sin \theta = \frac{ 3 }{ 5 }, \frac{ \pi }{ 2}\le \theta \le \pi \]

OpenStudy (anonymous):

 u have been visited by th ~dog of good fortune~ good fortune n small amounts of money will come 2 u v soon...........